OpenStudy (anonymous):

Explain in detail how to calculate the molar mass of K3PO4

OpenStudy (chmvijay):

add atomic mass in their stichiometric ratio example CO2 C=12 O=16 CO2 =12+2*16 =12+32=44

OpenStudy (anonymous):

K3=117 P=31 O4=64 molar mass=117+31+64=212